OpenStudy (anonymous):

Find the Taylor series for f(x)=1/(sqrt(x)) about x=9

OpenStudy (turingtest):

\[\sum_{n=0}^{\infty}{f^{(n)}(a)\over n!}(x-a)^n\]\[f(x)=x^{-\frac12}\implies f^{(n)}(x)=(-\frac12)^nx^{\frac12-n}\]\[\therefore \sum_{n=0}^{\infty}{f^{(n)}(9)\over n!}(x-9)^n={(-\frac12)^n9^{\frac12-n}\over n!}(x-9)^n\]

OpenStudy (turingtest):

I could write it in some slightly different ways of course is there a part you don't understand?

OpenStudy (turingtest):

slight typo in the exponent*\[\sum_{n=0}^{\infty}{f^{(n)}(a)\over n!}(x-a)^n\]\[f(x)=x^{-\frac12}\implies f^{(n)}(x)=(-\frac12)^nx^{-\frac12-n}\]\[\therefore \sum_{n=0}^{\infty}{f^{(n)}(9)\over n!}(x-9)^n={(-\frac12)^n9^{-\frac12-n}\over n!}(x-9)^n\]

OpenStudy (anonymous):

No not at all this is perfect I just had an answer and you have confirmed it. I've forgotten most things about Taylor series and I was just getting confirmation I was dong it correctly . Thanks very much

OpenStudy (turingtest):

very welcome :D